OpenStudy (anonymous):

Use implicit differentiation to find an equation of the tangent line to the devil's curve defined by y^2(y^2-9) = x^2(x^2-3^{2}+1) at the point (0,-3).

OpenStudy (anonymous):

So the equation is \[y^2(y^2-9)=x^2(x^2-3^2+1)\] ?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

OK first we are going to simplify it to make this easier for our sake. Distribute what we can into the parenthesis and also square the 3 to a 9 and we get. \[y^4 - 9y^2 = x^4 - 9x^2 + x^2\] You see how I got that?

OpenStudy (anonymous):

Mhmm I do

OpenStudy (anonymous):

Ok now here is the bugger. The derivative. Anytime we take a derivative of a term with y we need to tag it with dy/dx so lets do this a piece at a time \[\frac{ d }{ dx } y^4 = 4y^3\frac{ dy }{ dx }\] \[\frac{ d }{ dx } 9y^2 = 18y\frac{ dy }{ dx }\] \[\frac{ d }{ dx }x^4 = 4x^3\] \[\frac{ d }{ dx }9x^2 = 18x\] \[\frac{ d }{ dx }x^2 = 2x\] So then we have \[4y^3\frac{ dy }{ dx } - 18y \frac{ dy }{ dx }=4x^3 -16x\] This all make sense?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

now we solve to dy/dx by factoring it out and dividing by what is left \[\frac{ dy }{ dx }(4y^3-18y) = 4x^3-16x\] \[\frac{ dy }{ dx } = \frac{ 4x^3-16x }{ 4y^3-18x }\]

OpenStudy (anonymous):

This is our derivative now so now we got a few more steps to find the tangent line

OpenStudy (anonymous):

Okay! Thanks for your help so far :)

OpenStudy (anonymous):

Ok the point tells us that x = 0 and y = -3 so we plug those into our derivative to find the slope value at that point but as we can see if you plug in a 0 for x you get 0 in the numerator meaning our slope value is 0 which makes our job easier Set it up in slope intercept form y=mx+b y=0x+b y+b since y = -3 y=-3 is our tangent line

OpenStudy (anonymous):

y=b not y+b sorry

OpenStudy (anonymous):

So if we analyze this further it means that at (0,-3) This graph must have a horizontal tangent line to it